OpenStudy (anonymous):

How do I set up modelling equations for problems? For example Exponential growth. If the growth rate of the number of bacteria at any time t is proportional to the number present at t and doubles in 1 week, how many bacteria can be expected after 2 weeks? After 4 weeks?

OpenStudy (phi):

let x be the number of bacteria. The growth rate is the change in x over time \( \frac{dx}{dt}\) growth rate proportional to the number present means \[ \frac{dx}{dt} = k \cdot x \] where k is the constant of proportionality